# Permissive Information-Flow Analysis for Large Language Models

### v1

@@ V-C LLM agent dataset
TABLE III: Results on the LLM agent dataset. The utility target column indicates which output $y$ we are using to compute our target utility $U(p_{\textnormal{LM}}(y|x,C))$ in equation 1. In practice and in the absence of ground truth, the utility target is the model-generated output $y$, but we also compare with the ground truth $y^{\star}$ to understand the maximum possible utility assuming a perfect LLM. The difference in the numbers between the cases indicates that all the errors in the case of model-generated output are artifacts of the model generating a suboptimal response which causes incorrect propagation of labels.
@@
TABLE IV: Model output alignment with the dataset ground truth computed using ROUGE-L F-score on the LLM agent dataset. Utility computation is only applicable to the prompt-based label propagator. ROUGE-L$(y,y^{\star})$ refers to the alignment of the full context output (no regeneration required), while ROUGE-L$(y^{\prime},y^{\star})$ represents the alignment of the regenerated response after label improvement. Interestingly, we see that the regenerated response is more aligned with the ground truth response suggesting that in some cases label propagation can improve utility. ROUGE-L$(y^{\prime},y^{\star})-$ ROUGE-L$(y,y^{\star})$ compares the difference in alignment between the full context and the subcontext of the inferred label. When label improvement is possible, the difference in alignment is close to nil indicating that the regenerated response $y^{\prime}$ is highly aligned with the ground truth response $y^{\star}$. Otherwise, the difference in alignment is highly negative, indicating a significant drop in model alignment with $y^{\star}$.
In the LLM agent dataset, we take a step further by also computing the misalignment introduced by regenerating the output conditioned on the updated context (when label improvement is possible). The results are presented in Table III. We fix a threshold of $\lambda=0.2$ for this experiment. We use the base model without instruction tuning in this case due to the use of a custom chat format.
@@
To quantify this, we measure the alignment between $y$ and $y^{\star}$ as well as $y^{\prime}$ and $y^{\star}$ using the ROUGE-L F-score, with the results shown in Table IV. We obtain a ROUGE-L F-score of at least $0.85$, suggesting that conditioning on the reduced context leads to a similar model response. Furthermore, we see a negligible difference in alignment between the full context output and the reduced context output when label improvement is possible, but observe a significant drop when such an upgrade is not possible.
In a robust retrieval augmentation setup, adding additional documents to the context should not negatively influence the model’s ability to answer a query since the model is free to ignore irrelevant inputs [26]. In other words, we expect that the ground truth completion $y^{\star}$ is always similar to the completion given the full context $y$. However, in almost all label propagation errors we observe, we find that the model is not able to answer the query with the full context. When controlling the model generation by using the ground truth target in the dataset instead of using the model-generated output, we see that our label propagator achieves perfect precision and recall.
@@ VI Analysis and Discussion
RQ1. We find that our prompt-based label propagator is able to find the exact set of minimal labels in 86% of the cases. This is for a large lattice of labels where the label propagator needs to find the correct subset out of 16k possible sets of labels.
@@
RQ3. We showcase the label propagator in a real-world use case in a tool calling LLM agent setup where the propagated label is used to determine whether a sensitive tool call is allowed. We find that the RAG-based label propagator is able to improve the label in more than 85% of the cases while the output of the LLM agent remains the same as measured by the difference between the reduced context output alignment and the full context output alignment.
Comparison to baselines. The introspection-based label propagator achieves a noticeable improvement in performance when using a 70B parameter model compared to a 7B model. Furthermore, despite its simplicity and computational convenience, the introspection-based label propagator constitutes a strong baseline when considering simple lattices. We find a similar performance of the RAG-based label propagator for both model sizes suggesting that our approach could be implemented on smaller models, saving computational resources.
VI-A Use-cases
Algorithm 1 identifies sub-contexts with labels that are more permissive than that of the full context. However, whenever the desired output label $L$ can be determined up-front, it is possible to side-step the search over $\lambda$-similar sub-contexts and directly restrict the retrieval component to documents at or below $L$.
Our approach reveals its true benefits when the use of the generated content is not yet determined after the initial retrieval step. In such cases, having a too-restrictive label comes at the cost of limiting future uses of the generated output. For example:
  • Semantic caches [37] can be extended to store answers along with a sensitivity label, where more permissive labels facilitate broader reuse of the generated content.
  • Emails are often forwarded beyond their initial set of recipients, which is facilitated by using permissive labels.
  • Frameworks such as LangChain, LlamaIndex, Orkes, AutoGen, or TaskWeaver enable writing programs with LLM components. Similar to classic language-based information flow analysis [16], more permissive labeling enables the design of more expressive secure programs.
Lastly, our approach naturally allows users to endorse information. Let’s consider a setup as illustrated in Figure 1. Initially, there might only be an untrusted source answering the initial query. The LP system would correctly output an untrusted label to the potentially dangerous command. However, after a trusted authority (e.g., a university department IT admin) confirms the suggestion in the untrusted source, the LP recognizes that both sources are similarly influential and assign a trusted label to the output. Therefore, by quoting or repeating untrusted information a trusted source can endorse the information.
VI-B Limitations
Our label propagator is able to improve on the baseline label by more than 50% and 85% in two realistic datasets. However, this improvement comes with the cost of an increased number of LLM calls (discussed in Section III-E). We now highlight other limitations associated with our label propagator.
Impact of adversary
@@ VI-C Extensions
VII Related Work
Influence estimation
Influence estimation plays a crucial role in identifying the impact of individual training instances on a model’s predictions. Brophy et al. [39] present a good overview of different influence estimation techniques such as (i) leave-one-out estimation, (ii) Shapley value estimation, and (iii) sub-sample-based Shapley value estimation (adapted from Feldman and Zhang [40]). Influence estimation techniques have been popularly used in the machine learning literature for different applications. Koh and Liang [41] demonstrate the application of influence functions to debugging models and detecting dataset errors. Feldman and Zhang [40] explained the memorization of tail data points using the influence of those data points on the model’s predictions. Ilyas et al. [42] used influence estimation to present the idea of datamodels that allows the model to make predictions using linear regression. Cohen-Wang et al. [43] extended the use of datamodels to attribute the language model’s prediction to different parts of the model’s context window. Nguyen and Wong [44] presented a Shapley-value-based influence estimation technique to decide the best examples for in-context learning. Shapley value and its variants have been used to assign value to data contributed in collaborative data markets [45, 46, 47].
@@ VII Related Work
Information-flow analysis and the use of flow labels has a long-standing history, see e.g., [23], with an early survey of language-based approaches in [16]. The operations considered in many program analysis frameworks (e.g. [48, 16]) have well-defined semantics, and the rules for propagating the flow labels of each operator can be hard-coded. The situation is different when using ML models such as Transformers: due to their attention mechanism and autoregressive decoding, their output is tainted by all inputs. Moreover, their behavior is only implicitly defined as the solution of a loss-minimization problem. In our approach, we use this fact by propagating labels through ML components under the constraint that the model’s loss does not increase.
@@
Secure multi-execution [50, 51] is an approach to dynamic information-flow analysis where programs are executed once per security level. The idea is that an execution that produces output for level $L$ receives data only from levels $L^{\prime}\sqsubseteq L$; and dummy values from levels $L^{\prime\prime}\sqsubseteq L$. Programs that are multi-executed are secure by design, and correct if the original program was secure.
@@
The inscrutability of how AI components such as LLMs handle the information they are given makes it challenging to integrate them into larger systems while preserving confidentiality and integrity guarantees. A number of authors have explored how to provide best-effort or hard guarantees using techniques inspired by the information-flow analysis literature. Wutschitz et al. [52] study how to leverage existing metadata in datastores used for retrieval augmentation to enforce privacy guarantees at inference time, though they assume the target label of the output is given. Related to the introspection baseline we study, Mireshghallah et al. [53] use Contextual Integrity theory to explore how well LLMs can be relied upon to discern which information is appropriate to use in a given context. Similarly, Ghalebikesabi et al. [54] explore several ways to operationalize Contextual Integrity and find that using introspection gives promising results in a case study employing LLMs to fill in forms based on sensitive data. Finally, Wallace et al. [55] train LLMs to distinguish between instructions in their context totally ordered in a hierarchy, teaching them to prioritize higher-privileged instructions over instructions that appear lower in the hierarchy.
VIII Conclusion
@@
We presented a permissive approach to propagating information-flow labels of documents retrieved in RAG systems. The key idea is to propagate only the labels of those documents that are actually used for generating the model’s output. We show that our approach is practical in terms of performance and infers more permissive labels than an introspection baseline. Unlike introspection-based methods commonly used in practice, our approach can satisfy a hard safety guarantee.
We hope that this research spurs interest in system-level solutions to security and privacy concerns stemming from the use of unreliable AI software components. We may not be able to fully analyze how information flows through these components, but we can control which information they use and build more useful systems around them that enforce security and privacy policies.
IX Acknowledgements
The authors would like to thank colleagues at Microsoft for their helpful discussions and feedback.

### v2

@@ 5.3 LLM agent dataset
Table 3: Results on the LLM agent dataset. The utility target column indicates which output $y$ we are using to compute our target utility $U(p_{\textnormal{LM}}(y|x,C))$ in equation 1. In practice and in the absence of ground truth, the utility target is the model-generated output $y$, but we also compare with the ground truth $y^{\star}$ to understand the maximum possible utility assuming a perfect LLM. The difference in the numbers between the cases indicates that all the errors in the case of model-generated output are artifacts of the model generating a suboptimal response which causes incorrect propagation of labels.
@@
Table 4: Model output alignment with the dataset ground truth computed using ROUGE-L F-score on the LLM agent dataset. Utility computation is only applicable to the prompt-based label propagator. ROUGE-L$(y,y^{\star})$ refers to the alignment of the full context output (no regeneration required), while ROUGE-L$(y^{\prime},y^{\star})$ represents the alignment of the regenerated response after label improvement. Interestingly, we see that the regenerated response is more aligned with the ground truth response suggesting that in some cases label propagation can improve utility. ROUGE-L$(y^{\prime},y^{\star})-$ ROUGE-L$(y,y^{\star})$ compares the difference in alignment between the full context and the subcontext of the inferred label. When label improvement is possible, the difference in alignment is close to nil indicating that the regenerated response $y^{\prime}$ is highly aligned with the ground truth response $y^{\star}$. Otherwise, the difference in alignment is highly negative, indicating a significant drop in model alignment with $y^{\star}$.
In the LLM agent dataset, we take a step further by also computing the misalignment introduced by regenerating the output conditioned on the updated context (when label improvement is possible). The results are presented in Table 3. We fix a threshold of $\lambda=0.2$ for this experiment. We use the base model without instruction tuning in this case due to the use of a custom chat format.
@@
To quantify this, we measure the alignment between $y$ and $y^{\star}$ as well as $y^{\prime}$ and $y^{\star}$ using the ROUGE-L F-score (see Table 4). We obtain a ROUGE-L F-score of at least $0.85$, suggesting that conditioning on the reduced context leads to a similar model response. We see a negligible difference in alignment between the full context output and the reduced context output when label improvement is possible, but observe a significant drop when such an upgrade is not possible.
In a robust retrieval augmentation setup, adding additional documents to the context should not negatively influence the model’s ability to answer a query since the model is free to ignore irrelevant inputs [39]. Thus, we expect that the ground truth completion $y^{\star}$ is always similar to the completion given the full context $y$. However, in almost all label propagation errors we observe, we find that the model is not able to answer the query with the full context. When controlling the model generation by using the ground truth target in the dataset instead of using the model-generated output, we see that our label propagator achieves perfect precision and recall.
@@ 6 Analysis and Discussion
RQ1. We find that our prompt-based label propagator can find the exact set of minimal labels in 86% of the cases, for a large lattice of 16k possible labels.
@@
RQ3. We showcase the label propagator in a real-world use case in a tool calling LLM agent setup where the propagated label is used to determine whether a sensitive tool call is allowed. We find that the prompt-based label propagator is able to improve the label in more than 85% of the cases while the output of the LLM agent remains the same as measured by the difference between the reduced context output alignment and the full context output alignment.
Comparison to baselines The introspection-based label propagator achieves a noticeable improvement in performance when using a 70B parameter model compared to a 7B model. Furthermore, despite its simplicity and computational convenience, the introspection-based label propagator constitutes a strong baseline when considering simple lattices. We find a similar performance of the prompt-based label propagator for both model sizes suggesting that our approach could be implemented on smaller models, saving computational resources.
6.1 Use-cases
Algorithm 1 identifies sub-contexts with labels that are more permissive than that of the full context. However, whenever the desired output label $L$ can be determined up-front, it is possible to skip the search over $\lambda$-similar sub-contexts and restrict the retrieval component to documents at or below $L$.
Our approach reveals its true benefits when the use of the generated content is not yet determined after the initial retrieval step. In such cases, having a too-restrictive label comes at the cost of limiting future uses of the generated output. For example, semantic caches [3] can be extended to store answers along with a sensitivity label, where more permissive labels facilitate broader reuse of the generated content. Likewise, Emails are often forwarded beyond their initial set of recipients, which is facilitated by using permissive labels.
Our approach naturally allows users to endorse information. Let’s consider a setup as illustrated in Figure 1. Initially, there might only be an untrusted source answering the initial query. The LP system would correctly output an untrusted label to the potentially dangerous command. However, after a trusted authority (e.g., a university department IT admin) confirms the suggestion in the untrusted source, the LP recognizes that both sources are similarly influential and assigns a trusted label to the output. Therefore, by quoting or repeating untrusted information a trusted source can endorse information.
6.2 Limitations
Our label propagator is able to improve on the baseline label by more than 50% and 85% in two realistic datasets. However, this improvement comes with the cost of an increased number of LLM calls (discussed in Section 3.5).
@@ 6.3 Extensions
Efficiency improvements to Algorithm 1
Section 3.5 discusses the computational cost incurred to execute our label propagator in Algorithm 1. Despite an increase in the number of LLM calls for label propagation, the cost incurred per call can be drastically reduced utilizing common inference optimizations implemented in model serving backends [53, 21]. For instance, compute-bound prompt processing in Transformers can be amortized across calls sharing a common prompt prefix by reusing a KV cache. For a totally ordered lattice, appending retrieved documents at the end of the prompt sorted by their labels will result in no additional prompt processing costs in subsequent LLM calls because their prompts are obtained by peeling off documents at the end of the prompt.
We can also optimize the decoding phase of an LLM call by keeping a running calculation of the cumulative likelihood of the tokens decoded and stop decoding when the utility drops below a difference $\lambda$ of the utility of the full context.
Advanced influence estimation
The label propagation mechanism developed in Section 3 is based on estimating the influence of context elements on the model’s predictions. There are several alternatives techniques, such as Shapley value estimation, that can be used as drop-in replacements. See [4] for an overview of influence estimation techniques and [31, 8] for examples of their application to LLM contexts.
Another approach is to leverage the attention weights for influence estimation. Phenomena such as attention sinks [50] or lost-in-the-middle [25] makes identification based on simple attention scores difficult.
7 Related Work
@@
Information-flow analysis and the use of flow labels has a long-standing history, see e.g., [9], with an early survey of language-based approaches [35]. The operations considered in many program analysis frameworks (e.g. [44, 35]) have well-defined semantics, and the rules for propagating the flow labels of each operator can be hard-coded. The situation is different when using ML models such as Transformers: due to their attention mechanism and autoregressive decoding, their output is tainted by all inputs. Moreover, their behavior is only implicitly defined as the solution of a loss-minimization problem. In our approach, we propagate labels through ML components under the constraint that the model’s loss does not increase.
@@
Secure multi-execution (SME) [10, 34] is an approach to dynamic information-flow analysis where programs are executed once per security level. The idea is that an execution that produces output for level $L$ receives data only from levels $L^{\prime}\sqsubseteq L$; and dummy values from levels $L^{\prime\prime}\sqsubseteq L$. Programs that are multi-executed are secure by design, and correct if the original program was secure. Our approach is related to SME in that we also call the LLM once per security level, and we propagate the most permissive label for which the original output is still "correct", in the sense that the drop in utility is bounded by $\lambda$.
@@
Several emerging approaches rely on information-flow tracking to obtain deterministic security guarantees for LLM-based applications, even if the underlying models are vulnerable to attacks: Wutschitz et al. [48] study how to leverage existing metadata in datastores used for retrieval augmentation to enforce confidentiality guarantees at inference time. Wu et al. [47] propose a system that uses information-flow tracking to prevent LLM-based applications from being compromised by untrusted information. [2] develop a monitor and policy language that can check execution traces of LLM applications for violations of different safety and information-flow properties. In contrast to these approaches, our work aims to dynamically identify labels of information that is propagated through LLM calls, which can be used to increase the permissiveness of system-level information-flow tracking.
Related to the introspection baseline, Mireshghallah et al. [29] and Ghalebikesabi et al. [14] use Contextual Integrity theory to explore how well LLMs can be relied upon to discern which information is appropriate to use in a given context. Finally, Wallace et al. [45] train LLMs to distinguish between instructions in their context totally ordered in a hierarchy, teaching them to prioritize higher-privileged instructions over instructions that appear lower in the hierarchy.
8 Conclusion
@@
We presented a permissive approach to propagating information-flow labels of documents retrieved in RAG systems. The key idea is to propagate only the labels of those documents that are actually used for generating the model’s output. We show that our approach is practical in terms of performance and infers more permissive labels than an introspection baseline. Unlike commonly used introspection-based methods, our approach can satisfy a hard safety guarantee.
9 Open Science
We will open-source our prototype implementation and the datasets used for its evaluation.
10 Ethics Considerations
We use a public dataset of fake and real news articles [33] as well as fully synthetic datasets for our evaluations. For instance, the synthetic key-value dataset contains entirely generated Social Security Numbers (SSNs) and dates of birth (DoB). Therefore, no private information from any individual is collected or used. Similarly, the LLM agent dataset is entirely synthetic and generated by us, ensuring that no real user interactions were recorded.
Additionally, we present a label propagation technique that can positively impact users of LLMs by providing hard safety guarantees, enhancing their reliability and security.
Appendix A Improving performance on fine-grained lattices
Figure 9: Hyperparameter grid search on Shapley threshold and perplexity tolerance for the 7B (left) and 70B (right) models on the synthetic key-value dataset.
Figure 10: Hyperparameter grid search on Shapley threshold and perplexity tolerance for the 7B (left) and 70B (right) model on the news article dataset with a perfect retriever (first row) and realistic retriever (second row).
Algorithm 1 relies only on $\lambda$ to produce the minimal set of labels. However, in the cases where documents share similarities and provide important information regarding the model output such as the format (which is the case for our synthetic key-value dataset 4.5), the overall contribution of irrelevant documents goes up. Therefore, additional irrelevant documents are introduced in the predicted label set due to being $\lambda$-similar.
In order to cover these cases when dealing with a complex lattice and a large number of relations, we propose a simple heuristic based on Shapley values that have been commonly used in the past [37, 31]. In particular, we compute the Shapley value for each of the different labels in the lattice and filter out any label combinations where labels below a particular threshold are present.
The Shapley value defines the marginal contribution of any label $L$ by computing the average difference in outcomes when a particular label is present and absent. We compute the label Shapley value via perplexity, which is computed as the average drop in perplexity when a particular label is included. This provides a notion of the importance of each label. This now introduces an additional hyperparameter i.e., the Shapley value threshold.
We visualize the results for the 2D grid search by considering both perplexity tolerance $\lambda$ as well as Shapley threshold on the synthetic key-value dataset in Figure 9, and the news article dataset in Figure 10.
The results indicate that on simpler lattices such as in the case of news articles, we see almost no impact due to the Shapley value threshold. On the other hand, for complex lattices, such as in the case of our synthetic key-value dataset, we observe 6% absolute improvement in the exact-match accuracy, highlighting the utility of this heuristic in such cases.